OpenStudy (anonymous):

What is the acceleration due to gravity at an altitude of 1.00 × 106 m above the earth's surface? Note: the radius of the earth is 6.38 × 106 m.

OpenStudy (anonymous):

I have been on this problem for 30 min

OpenStudy (mendicant_bias):

No worries, bud. So, I'm guessing in that second line, the altitude is \[1.0\times10^{6}\] yes?

OpenStudy (anonymous):

yes

OpenStudy (mendicant_bias):

Alright. Mind me asking what formulas you think might be relevant here? Just throw out some ideas.

OpenStudy (anonymous):

i tried to use g=GM/R^2

OpenStudy (mendicant_bias):

Alright, good! That's one, but you'll need one more. You'll need to get two equations involved, and use some substitution or equating in order to eliminate some things. Got an idea what the other might be?

OpenStudy (mendicant_bias):

I'll give you a hint: It's the one Newton is easily most famous for.

OpenStudy (anonymous):

E=mc^2 sorry I'm lost

OpenStudy (mendicant_bias):

Oh, and, nevermind, actually, you had the right one. I misread, if you only have that, that's all you need. Lol, no problem.

OpenStudy (mendicant_bias):

I'm going to take a little bit of a different approach and assume that you don't yet have F = GM/r^2, to make a point. Let's start out with some universal or given things, yes?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use the formula g=GM/r^2

OpenStudy (anonymous):

f is not equal to GM/r^2

OpenStudy (anonymous):

f=Gm1m2/r^2

OpenStudy (mendicant_bias):

You have Newton's 2nd Law, which, when written, appears as: \[F = ma.\]This works for all objects which have a constant mass, so this should be everything you are dealing with for now. In this question, the object you are dealing with is presumably only being acted on by gravity, so you can reasonably say that \[F_{g} = mg\]Where g is the gravitational acceleration caused by Earth pulling on your object. Next, you need to get Newton's Universal Lawof Gravitation involved. \[F = \frac {GmM}{R^{2}}\] You can straightforward plug in Newton's 2nd Law of Motion into his Law of Universal Gravitation, cancel one of the masses, and be left with the equation you started with. That's a good starting point to get out of the way in the first place. That's where that F = Gm/R^2 came from in the first place, so you understand. Now, onto your problem.

OpenStudy (mendicant_bias):

How do you think you can find the total radius, R, of the object that is some altitude above the Earth's surface? What is it?

OpenStudy (anonymous):

6400km + given distance

OpenStudy (mendicant_bias):

(Answering the question for people generally isn't how people learn, but thanks...)

OpenStudy (anonymous):

or 6.38x106

OpenStudy (anonymous):

Rtotal= radius of earth plus altitude, so i did 1.00 10^6 + 6.38 10^6

OpenStudy (anonymous):

thats true. i just wanted to correct the equation

OpenStudy (anonymous):

mendicant..have you done rocket propulsion?

OpenStudy (mendicant_bias):

A little bit; I'm gonna try to take a shot at yours in a sec.

OpenStudy (anonymous):

kk thx

OpenStudy (mendicant_bias):

So, now, you have GM/r^2, and should have values for G, M, and R, assuming you have been given (or can look up) the Gravitational Constant.\[F = \frac {GmM}{R^{2}}\]\[F_{g} = mg\]\[mg = \frac {GmM}{r^{2}} = \frac {GM}{R^{2}}\]--- \[g = \frac {GM}{R^{2}} \]

OpenStudy (mendicant_bias):

From there just plug and solve, yes?

OpenStudy (anonymous):

but is R the radius + altitude

OpenStudy (mendicant_bias):

Physo, you misunderstand. It *is* equal to GM/R^2 in this instance. You posted Newton's Universal Law of Gravitation, which is definitely relevant, but is not the exact same ting. Yes, R is the radius plus altitude. That's what R is. What's the problem in solving?

OpenStudy (anonymous):

I think i might just be entering it into the calculator wrong, Thank You very much for really taking the time to help me

OpenStudy (mendicant_bias):

No problem! You already had the solution anyways, it seems. Let me take a shot to make sure we're getting the same thing.

OpenStudy (anonymous):

the accelleration for all things falling or going upwards on earth is always either 9.8 m/s/s or -9.8 m/s/s

OpenStudy (anonymous):

it actually differ by a tiny amount with altitude.

OpenStudy (anonymous):

differs*

OpenStudy (anonymous):

oh I never learned that!

OpenStudy (mendicant_bias):

\[F_g = \frac {GM}{R^{2}} = \frac {(6.67\times10^{-11})(5.97\times10^{24})}{(6.38\times10^{6}+1.00\times10^{6})^{2}} = \frac {(6.67\times10^{-11})(5.97\times10^{24})}{(7.38\times10^{6})^{2}}= \] (Sorry, wrting it out tediously and slowly as possible)

OpenStudy (anonymous):

so it was the notation which caused all the misunderstanding. By Fg, you mean acceleration due to gravity

OpenStudy (anonymous):

right?

OpenStudy (mendicant_bias):

\[\frac {3.98\times10^{14}}{5.34\times10^{13}} = 7.31 \ \text {m/s}\] Man, LTeX is just not playing nice tonight There we go.Did you get that?

OpenStudy (mendicant_bias):

And yeah, just a notation difference. *LaTeX.

OpenStudy (anonymous):

Yes i did !!!!!!!

OpenStudy (anonymous):

Thank You

OpenStudy (mendicant_bias):

No problem, man! You're welcome. You already knew how to solve it, heh, just human error.